I need help on 3/5(x+2)=12
Answer by checkley77(12844)

3/5(x+2)=12
3x/5+3/5*2=12
3x/5+6/5=12
(3x+6)/5=12
3x+6=5*12
3x+6=60
3x=60-6
3x=54
x=54/3
x=18 ans.
Proof:
3/5(18+2)=12
3/5*20=12
60/5=12
12=12